OpenStudy (anonymous):

Given b = 2 and h = 8, what is the equation of the graph if the parent function is y equals the square root of x? A. y= sqrt(2x+16) B. y= sqrt(2x-16) C. y= sqrt(2x) +16 D. y= sqrt(2x) -16

OpenStudy (anonymous):

This question does not have any direction about what to do with b or h. Could you provide more information?

OpenStudy (anonymous):

b=2 h-8 but what gets me stuck is that 16 is where h is? but b is in the right place its supposed to be like plugging in the values but h isn't even there? 16 is

OpenStudy (anonymous):

@jkl5149

OpenStudy (anonymous):

h=8*

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

I'm still unsure about what to do with the problem. I don't know why b or h would be there, is there any other information about the problem you aren't telling me?

OpenStudy (anonymous):

no that's all that's on the problem...

OpenStudy (anonymous):

Hmm, sorry then I can't help you. I don't see where the b or h would go. :(

OpenStudy (anonymous):

do you have anyone that could help?

OpenStudy (anonymous):

Ỏ̷͖͈̞̩͎̻̫̫̜͉̠̫͕̭̭̫̫̹̗̹͈̼̠̖͍͚̥͈̮̼͕̠̤̯̻̥̬̗̼̳̤̳̬̪̹͚̞̼̠͕̼̠̦͚̫͔̯̹͉͉̘͎͕̼̣̝͙̱̟̹̩̟̳̦̭͉̮̖̭̣̣̞̙̗̜̺̭̻̥͚͙̝̦̲̱͉͖͉̰̦͎̫̣̼͎͍̠̮͓̹̹͉̤̰̗̙͕͇͔̱͕̭͈̳̗̭͔̘̖̺̮̜̠͖̘͓̳͕̟̠̱̫̤͓͔̘̰̲͙͍͇̙͎̣̼̗̖͙̯͉̠̟͈͍͕̪͓̝̩̦̖̹̼̠̘̮͚̟͉̺̜͍͓̯̳̱̻͕̣̳͉̻̭̭̱͍̪̩̭̺͕̺̼̥̪͖̦̟͎̻̰_Ỏ̷͖͈̞̩͎̻̫̫̜͉̠̫͕̭̭̫̫̹̗̹͈̼̠̖͍͚̥͈̮̼͕̠̤̯̻̥̬̗̼̳̤̳̬̪̹͚̞̼̠͕̼̠̦͚̫͔̯̹͉͉̘͎͕̼̣̝͙̱̟̹̩̟̳̦̭͉̮̖̭̣̣̞̙̗̜̺̭̻̥͚͙̝̦̲̱͉͖͉̰̦͎̫̣̼͎͍̠̮͓̹̹͉̤̰̗̙͕͇͔̱͕̭͈̳̗̭͔̘̖̺̮̜̠͖̘͓̳͕̟̠̱̫̤͓͔̘̰̲͙͍͇̙͎̣̼̗̖͙̯͉̠̟͈͍͕̪͓̝̩̦̖̹̼̠̘̮͚̟͉̺̜͍͓̯̳̱̻͕̣̳͉̻̭̭̱͍̪̩̭̺͕̺̼̥̪͖̦̟͎̻̰

OpenStudy (irishboy123):

you can massage the numbers if you look at it thus: \(y = \sqrt{ b(x+h) } \) or \(y = \sqrt{ b(x-h) } \) but, if i understand this thread correctly, the right thing i think is to seek further clarification as you're answering a question that you have to guess. that's a bit of a handicap.